OpenStudy (aundiewear):

can some one help me with the sum of series?? I'm going to write the problem >>

OpenStudy (aundiewear):

\[\sum_{k=1}^{\infty}5(1/\sqrt{5})^k= ?\]

ganeshie8 (ganeshie8):

whats the first term in the sum ?

OpenStudy (unklerhaukus):

√5

OpenStudy (unklerhaukus):

what is the next term? (when k=2)

OpenStudy (aundiewear):

k=1